NASA International Space Station On-Orbit Status 4 May 2016

Scott Kelly uses ultrasound to image his jugular vein with the aid of Gennady Padalka, in order to evaluate the effectiveness of a lower body negative pressure countermeasure (Chibis LBNP, controlled by Mikhail Kornienko) used to reverse the headward fluid shift that occurs in the weightlessness environment of space. Credit: NASA.

The astronauts onboard the International Space Station are researching how microgravity affects fluid shifts in a crew member's body. Ground controllers are also guiding Canada's robotic arm into position before next week's grapple and release of the SpaceX Dragon spacecraft.

The Fluid Shifts experiment will wrap up operations this week with the crew wearing specialized body suits. The suits, known as Chibis Lower Body Negative Pressure devices, measure how fluids move from the lower body to the upper body while living in space. The research also observes fluid shifts in and out of cells and blood vessels which may impact head pressure potentially affecting vision.

Robotics controllers are remotely guiding the Canadarm2 to the Harmony module where it will grapple Dragon ahead of the spacecraft's May 11 release and splashdown in the Pacific Ocean. Yesterday, the controllers surveyed the Alpha Magnetic Spectrometer's condition with the Canadarm2 and its cameras.

Cosmonaut Oleg Skripochka is on his second mission aboard the space station. His first mission during Expedition 25/26 lasted 159 days. Currently, he is Expedition 47 Flight Engineer and today is his 47th day aboard the orbital lab since his March arrival with fellow crew members Jeff Williams and Alexey Ovchinin.

On-Orbit Status Reports

Fluid Shifts Baseline Imaging Using Chibis in the Russian Segment: The crew began operations for the third and final week of the Fluid Shifts experiment by performing data collections while wearing the Chibis Lower Body Negative Pressure (LBNP) device. The Fluids Shift investigation is divided into three one-week segments: Dilution Measures, Baseline Imaging, and Baseline Imaging using the Russian Chibis Lower Body Negative Pressure (LBNP) device. The experiment measures how much fluid shifts from the lower body to the upper body, in or out of cells and blood vessels, and determines the impact these shifts have on fluid pressure in the head, changes in vision and eye structures.

Rodent Research-3 (RR-3) Transfer and Bone Densitometer Scans: Following yesterday's grip strength operations, today the crew transferred the rodents from their habitats, configured Ultrasound 2, and performed bone densitometer scans with support from ground teams. This is one of three sessions planned this week that studies the molecular and physical changes to the musculoskeletal system that occur in space. Results will expand scientists' understanding of muscle atrophy and bone loss in space while testing an antibody that has been known to prevent muscle wasting in mice on Earth.

Microbial Observatory-1 Review: In preparation for upcoming Microbial Observatory -1 operations, the crew reviewed reference material and videos before discussing sample collection procedures with the payload developer. The Microbial Observatory-1 investigation monitors the types of microbes present on ISS over a one-year period. Samples are returned to Earth for further study, enabling scientists to understand the diversity of the microbial flora on the ISS and how it changes over time.

AMS external survey: Last night, Robotics Ground Controllers successfully completed a survey of the Alpha Magnetic Spectrometer (AMS) payload from various angles to understand the state of the payload as well as to obtain detailed views of the payload. This evening, controllers will translate the Mobile Transporter (MT) from Work Site 2 (WS2) to WS4. The Space Station Remote Manipulator System (SSRMS) will then walk off to Node 2 to perform Dragon Thermal Protection System (TPS) inspection and re-grapple in preparation for unberth on May 11.

Treadmill 2 (T2) software update: Ground teams updated the software on the T2. The regularly scheduled exercise session on T2 after the update was performed nominally.

Pressurized Mating Adapter (PMA1)/Node 1 Intermodule Ventilation (IMV) Reconfiguration: Following the PMA1/Node 1 IMV reconfiguration activity on Monday, data indicated that air exchange is still occurring between the Russian segment and the USOS. Today, to limit the air exchange through the PMA, the crew repositioned the Node 1 Overhead Starboard diffuser vanes so the air flows away from the PMA. Additionally, the crew will be reconfiguring the stowage in the PMA to direct the air coming out of the PMA grille toward the Russian Segment. The airflow between the segments is being limited in order to optimize CO2 removal throughout the ISS.

Today's Planned Activities
All activities were completed unless otherwise noted.

FLUID SHIFTS. Comm configuration for the experiment / r/g 9995
FLUID SHIFTS. Gathering and Connecting Equipment for TV coverage
FLUID SHIFTS. Activation of РБС for Ultrasound Equipment / r/g 2133
FLUID SHIFTS. Ultrasound 2 Setup and Activation in SM
Rodent Research (RR) Camcorder Setup
FLUID SHIFTS. Assistance with Chibis and Gamma-1 r/g 2136
FLUID SHIFTS. Chibis Setup / r/g 2136
Rodent Research (RR) Animal transfer to MSG and back
Rodent Research (RR) Blood Samples Retrieval from MELFI
Replacement of ПФ1, ПФ2 Dust Filter cassettes and cleaning В1, B2 fan screens in MRM2
Rodent Research (RR) Experiment Ops
Rodent Research (RR) Bone Densitometer Scan
FLUID SHIFTS. Ultrasound 2 Scan, Operator 1
FLUID SHIFTS. Ultrasound 2 Scan, Subject
Rodent Research (RR) Bone Densitometer Scan
СКПФ1, СКПФ2 Dust Filter Replacement and MRM1 Gas-Liquid Heat Exchanger (ГЖТ) Cleaning
FLUID SHIFTS. Chibis Closeout Ops / r/g 2136
Rodent Research (RR) Camcorder Setup
Gathering and connecting SM-AGAT-U52 charger r/g 2130
Verification of ИП-1 Flow Sensor Position
24-hour ECG Monitoring (start) / r/g 2107
24-hour BP monitoring (start) / r/g 2108
MICRO10 SABL 2 Rack Doors Open
Complete the Dose Tracker application
Rodent Research (RR) Habitat Transfer
FLUID SHIFTS. Chibis Setup / r/g 2135
FLUID SHIFTS. Operator Assistance with Chibis and Gamma-1 r/g 2135
FLUID SHIFTS. Ultrasound 2 Setup and Activation in SM
ПФ1, ПФ2 Dust Filter Replacement and В1, В2 Fan Grille Cleaning in DC1
FLUID SHIFTS. Ultrasound 2 Scan, Test Subject
FLUID SHIFTS. Ultrasound 2 Scan, Operator
Vacuum Cleaning of ВД1 and ВД2 air ducts in DC1
FLUID SHIFTS. Chibis Closeout Ops / r/g 2135
SPHEROIDS. Temperature Check
FLUID SHIFTS. Deactivation of РБС and USND / r/g 2134
Soyuz 720 Samsung Tablet Recharge, Initiate
Soyuz 719 Samsung Tablet Recharge, Initiate
Rodent Research (RR) Habitat Transfer
Rodent Research (RR) Experiment Ops
FLUID SHIFTS. Ultrasound 2 Stowage in SM
Rodent Research (RR) Bone Densitometer Scan
FLUID SHIFTS. CCFP hardware, HRF Laptop in SM
PMA1 airflow redirect [Deferred]
Rodent Research (RR) Bone Densitometer Scan
Cleaning of TVS LIV voltage converter vent grille
СОЖ Maintenance
VIZIR. Experiment Ops with СКПФУ Hardware r/g 2125
FLUID SHIFTS. Installation of cable assembly for voltage converter r/g 2131
FLUID SHIFTS. Hardware Handover to USOS
Cleaning FGB ЦВ1 Circulation Fan Screens
FSL Optical Adapter Cable removal and restoring EDR and FSL connections
Rodent Research (RR) Experiment Ops
FLUID SHIFTS. Connecting US-invertor to СУБА
FLUID SHIFTS. Crew Onboard Support System (КСПЭ) Hardware Deactivation and Closing Applications
FLUID SHIFTS. Restore nominal comm config
CONTENT. Experiment Ops / r/g 2119
USND2 Verification of devices and camcorder configuration
Rodent Research (RR) Habitat Transfer
Reconfigure LAB camcorder and Dual Canon Li-Ion Battery Charger to nominal config
MICROBS1 Procedure Review
CONTENT. Experiment Ops / r/g 2118
MICROBS1 Crew Conference
Soyuz 720 Samsung Tablet Recharge, Terminate
Flushing БКО for Condensate Water Recovery System r/g 2121
Soyuz 719 Samsung Tablet Recharge, Terminate
C2V2 Voice Check

Completed Task List Items
WHC KTO replace

Ground Activities
All activities were completed otherwise noted.
RR ops support
C2V2 checkout
Nominal ground commanding

Three-Day Look Ahead:
Thursday, 05/05: Fluid Shifts, Nano Racks Platform 2 module removal, Microbial sample collect
Friday, 05/06: RR CASIS ops, Nano Racks Module 9 ops, Fluid Shifts, Microbial sample collect
Saturday, 05/07: Crew off duty, housekeeping

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Idle
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) Standby
Urine Processing Assembly (UPA) Standby
Trace Contaminant Control System (TCCS) Lab Off
Trace Contaminant Control System (TCCS) Node 3 Full Up

Please follow SpaceRef on Twitter and Like us on Facebook.